OpenStudy (anonymous):

-3X+2 over 4 = 2X+1 over -3

OpenStudy (anonymous):

The easiest way to get rid of the division. Multiply everything by 4 to get rid of the division. \[-3x+2=4(2x+1)/-3\]Next, multiply everything by -3 to get rid of the division by -3. \[-3(-3x+2)=4(2x+1)\]Distribute. \[+9x-6=8x+4\]Now you've gotten rid of the division. You eventually want this thing to look like x=#, which means you want all the x's on one side and everything else on the other. Let's get rid of the 8x from the right side by subtracting it. That means we have to subtract 8x from the left side, too. \[1x+2=1\]We don't want any constants on the left side, so we add 6 to both sides. \[x=10\] Then check to make sure the answer works by plugging in 10 for x.

OpenStudy (anonymous):

I skipped a couple steps, so let me know if you have any questions about it.

OpenStudy (anonymous):

thank you !